# Output Range Analysis for Deep Neural Networks based on Simulated Annealing Processes

### v1

@@ 6 Experimental Evaluation
To illustrate the use and operation of our algorithm we are going to use the Ackley function. In mathematical optimization, the Ackley function is a non-convex function, with many local minima, which is often used as a performance test problem for optimization algorithms.This function on a 2-dimensional domain it is defined by
@@
for which its global minimum point is $f(0,0)=0$, see Figure 6.1. In this case we are going to assume that $E=[-4,4]\times[-4,4]$. Furthermore, suppose we want to estimate the function (14) using only a discretized sample of points $\mathcal{D}=\{(x_{1i},x_{2i},f_{i})\}_{i=1}^{m}\subset E$. To this end, we use a ResNet as described in Chapter 2, with $L=5$ and $H_{l}=10$ for $l=2,\dots,5$. The fit generated by these ResNet is quite good, the surface generated by the neural network maintains all the properties of the original function, including its non-convex nature and the existence of many local minima, See left side of Figure 6.2.
Figure 6.1: Ackley function of two variables
Figure 6.2: Left: Surface generated by the ResNet $(E,\mathcal{F}(E))$. Right:
@@
The fact that the generated surface consists of too many local minima constitutes a suitable scenario where our methodology shows its greatest strength. Let us remember that when working with deep neural networks, this scenario is quite recurrent due to the high non-linearity and flexibility of the relationships established by this type of neural networks. After training the ResNet $\mathcal{F}$, our main objective is to find the global minimum on the surface generated with the neural network within the domain $E$, denoted as $(E,\mathcal{F}(E))$. To this end, we apply our Algorithm 1 to this case. On the right side of Figure 6.2 we present the evolution of the values assumed by the Markov process generated through the algorithm. On the right side of Figure 6.2 we present the evolution of the values assumed by the Markov process generated through the algorithm. We note the reasonably fast convergence towards the global minimum $\mathcal{F}_{\min}=0$. While it is true that there are fluctuations around the trajectory towards convergence, this is due to the highly fluctuating nature of the generated surface $(E,\mathcal{F}(E))$, but there is also a clear and consistent trend towards the global minimum 00. The Python codes of this experimental evaluation are available in https://github.com/Nicerova7/output-range-analysis-for-deep-neural-networks-with-simulated-annealing.
@@ 7 Conclusions
In this paper, we presented a novel approach to the output range analysis problem for Deep Neural Networks (DNNs) by employing Simulated Annealing (SA). Our methodology was specifically designed to handle the lack of local geometric information and the high non-linearity typical of DNNs. Through detailed theoretical analysis and extensive experimental evaluations, we demonstrated the effectiveness of our algorithm across a wide spectrum of neural networks, with a particular focus on Residual Networks (ResNets) due to their prominence in applications.
We began by mathematically describing ResNets and defining the output range analysis problem, highlighting the challenges posed by the complex and non-linear nature of DNN response surfaces. We then adapted the SA algorithm to this context, ensuring it could work within the bounded domains typical of DNN inputs.

### v3

@@ 7.1 The Ackley Function
In mathematical optimization, the Ackley function is a non-convex function, with many local minima, which is often used as a performance test problem for optimization algorithms.This function on a 2-dimensional domain is defined by
@@
for which its global minimum point is $f(0,0)=0$, see Figure 1(a). We use this function to obtain a noisy discretized sample $\mathcal{D}=\{(x_{1i},x_{2i},f_{i})\}_{i=1}^{m}\subset E$, where $E=[-4,4]\times[-4,4]$, see 1(b). Furthermore, suppose we want to estimate the function (14) using only the discretized sample of points $\mathcal{D}$. To this end, we use a ResNet as described in the chapter “Residual Neural Networks,” with a total depth of $L=5$ residual blocks. The input layer maps a 2-dimensional input to $H_{0}=128$ neurons. The network contains 4 intermediate residual blocks, each with $H_{l}=256$ neurons for $l=1,\dots,4$, followed by a fifth residual block that reduces the dimensionality back to $H_{5}=128$ neurons before the output layer. The output layer maps the final $128$ neurons to a single output. All layers in the network use the ReLU activation function. The model was initialized with the following setup: Mean Squared Error (MSE) was employed as the loss function, and the Adam optimizer kingma2017adam was used smooth surface generated by ResNet for training with a learning rate of 0.001. The model was trained for 1000 epochs. The fit generated by these ResNet is quite good, resulting in a mean absolute error of $1.0070$ and a mean squared error of $3.0298$ from $1000$ uniform random points generated within the range $[-5,5]\times[-5,5]$. The surface generated by the neural network maintains all the properties of the original function, including its non-convex nature and the existence of many local minima, see Figure 1(c). As mentioned above, we will use this generated response surface to find the global minimum of the Ackley function.
(a) Ackley function in two dimensions.
(b) Discretized sample from the Ackley function
(c) Response surface generated (smoothing) from the ResNet
Figure 7.1: Graphs of the experimental evaluation using the Ackley function.
@@
The fact that the generated surface consists of too many local minima constitutes a suitable scenario where our methodology shows its greatest strength. Let us remember that when working with deep neural networks, this scenario is quite recurrent due to the high non-linearity and flexibility of the relationships established by this type of neural network. After training the ResNet $\mathcal{F}$, our main objective is to find the global minimum on the surface generated with the neural network within the domain $E$, denoted as $(E,\mathcal{F}(E))$. To this end, we apply our Algorithm  1 to this case. In figure 7.2, we present the evolution of the values assumed by the Markov process generated through the algorithm and its tendency. We note the reasonably fast convergence towards the global minimum $\mathcal{F}_{\min}=0$. While it is true that there are fluctuations around the trajectory towards convergence, this is due to the highly fluctuating nature of the generated surface $(E,\mathcal{F}(E))$, but there is also a clear and consistent trend towards the global minimum 00.
Figure 7.2: Red line: Sample path of Markov process $(X_{i})_{i\in\mathbb{N}}$ for the Ackley function. Blue line: fitted trend line.
7.1.1 Differences with the Classical Simulated Annealing Algorithm
As mentioned above, the fundamental difference between our proposed algorithm and the classical Simulated Annealing algorithm is the incorporation of a restricted search domain. This domain restriction is a priority need present in applications within the area of output range analysis for neural networks. As shown in Section 5, we successfully incorporate bounded domains into the algorithm and further establish theoretical properties that guarantee the convergence of the algorithm to optimal points. In this subsection, ignoring the priority need for domain restrictions, for performance comparison purposes only, we apply the classical SA algorithm with a scheme analogous to Algorithm 1 but without the incorporation of bounded domains. In Figure 7.3 shows the extensive range and unbounded path traced by the classic Simulated Annealing algorithm. In iteration 125 a maximum value, $131.21$, is found to subsequently fluctuate to converge to the minimum. Therefore, the non-incorporation of a restricted domain causes the algorithm to experience large fluctuations before converging to the optimal points. But, as we mentioned before, this comparison is only didactic, since restricted domains are part of the nature of the problem. In many cases, as part of the diagnosis of the trained neural network, it is necessary to evaluate the optimal points assumed by the neural network in certain previously defined limited domains.
Figure 7.3: Red line: Sample path of the classical Simulated Annealing algorithm for the Ackley Function. Blue line: fitted trend line.
7.2 The Drop-Wave Function
This function is characterized by being multimodal, with many local minima, and highly complex; it is also non-convex. Like the previous function, it has a 2-dimensional domain and is defined by
$f(x_{1},x_{2})=-\frac{1+\cos\left(12\sqrt{x_{1}^{2}+x_{2}^{2}}\right)}{0.5(x_{1}^{2}+x_{2}^{2})+2},$ (15)
for which its global minimum point is $f(0,0)=-1$, see Figure 7.4. In particular, the function is usually evaluated on the square $E=[-5.12,5.12]\times[-5.12,5.12]$. Moreover, suppose we want to estimate the function (15) using only a discretized sample of points $\mathcal{D}=\{(x_{1i},x_{2i},f_{i})\}_{i=1}^{m}\subset E$. For this experiment, we use a total depth of $L=7$ residual blocks. The input layer maps a 2-dimensional input to $H_{0}=128$ neurons. The network contains $L=7$ residual blocks with the following dimensions: $H_{1}=H_{2}=256$, $H_{3}=H_{4}=H_{5}=512$, $H_{6}=256$, and $H_{7}=128$. Finally, The output layer has $128$ neurons. All layers in the network use the ReLU activation function The model was configured with the following setup: The loss function utilized was Mean Squared Error (MSE). Training was performed using the Adam optimizer kingma2017adam with a learning rate of $0.001$ and was performed over 1000 epochs. The fit generated by these ResNets is strong, with a mean absolute error of $0.0221$ and a mean squared error of $0.0010$, based on 1500 points generated within the range $[-5.12,5.12]\times[-5.12,5.12]$. The surface generated by the neural network maintains all the properties of the original function, including its non-convex nature and the existence of many local minima, see Figure 7.5.
Figure 7.4: Drop-Wave function of two variables. Left: Perspective view; Right: Top View.
Figure 7.5: Drop-Wave surface generated by the ResNet $(E,\mathcal{F}(E))$. Left: Perspective view; Right: Top View.
Figure 7.6: Red line: Sample path of Markov process $(X_{i})_{i\in\mathbb{N}}$ for the Drop-Wave function. Blue line: fitted trend line.
This second function is highly complex, and its experiment demonstrates the power of our method. Despite the presence of numerous local minima, our method successfully approximates to the global minimum. This scenario is crucial for solving highly complex problems, which are common in physics and other fields Bezhko_2021; physic_plasma. We applied our Algorithm  1 and obtained the results shown in Figure 7.6. The evolution of values assumed by the Markov process in this instance reveals the full path, where the low values we seek do not always appear at the end of the iteration, but they are still identified. The global minimum, $\mathcal{F}_{\min}=-1$, was successfully found using our method. The trend line clearly illustrates the fluctuations, making them easier to observe. Similar to the previous experiment with the Ackley function, the high fluctuations in the generated surface explain the highly fluctuating path we observed.
7.2.1 Differences with the Classical Simulated Annealing Algorithm
As in the previous case, we forget for a moment the imperative need for restricted domains. Therefore, only for diagnostic purposes will we compare our proposal with the classic SA. In this highly complex example, we found an acceptable mapping DNN for The Drop-Wave function (15). In Figure 7.7, we observed a similar behavior between this classic SA algorithm and our method until iteration $300$. However, the chaotic movement of the classic SA shifted significantly towards the opposite direction of the minimum value. Moreover, if we examine closely, these values do not make sense for the original equation (15), which is easily distinguishable in Figure 7.5.
Figure 7.7: Red line: Sample path of the classical Simulated Annealing algorithm for the Drop-Wave function. Blue line: fitted trend line.
7.3 Multiple Global Minima Function
Our method also can be applied in an space where there are multiple global minima. Also, it is not restricted for two dimension but more than two dimensions, and in particular n-dimensions, is allowed. For example, consider the function
$f(x,y,z)$ $=(x^{2}-1)^{2}+(y^{2}-1)^{2}+(z^{2}-1)^{2},$ (16)
where each term $(x^{2}-1)^{2}$, $(y^{2}-1)^{2}$, and $(z^{2}-1)^{2}$ achieves its minimum value of 0 when $x^{2}=1$, $y^{2}=1$, and $z^{2}=1$, respectively. This results in two possible values for each variable: $x=\pm 1$, $y=\pm 1$, and $z=\pm 1$. The combination of these values gives the following 8 global minima:
[…]
. The function is evaluated on the square $E=[-3,3]\times[-3,3]\times[-3,3]$. Then, we want to estimate the function (16) using only a discretized sample of points $\mathcal{D}=\{(x_{1i},x_{2i},,x_{3i},f_{i})\}_{i=1}^{m}\subset E$.
Figure 7.8: Red line: Sample path of Markov process $(X_{i})_{i\in\mathbb{N}}$ for the Multiple Global Minima function. Blue line: fitted trend line.
For this function, we use $L=7$ residual blocks, where $H_{0}=128$ represents the neurons in the input layer, $H_{1}=H_{2}=256$, $H_{l}=512$ for $l=3,4,5$, $H_{6}=256$, and $H_{7}=128$, which represents the neurons in the final residual block. The output layer maps the final $128$ neurons to a single output. All layers in the network use the ReLU activation function, except for the final output layer. The model was configured with the following setup: The loss function utilized was Mean Squared Error (MSE). Training was performed using the Adam optimizer with a learning rate of $0.001$ and was performed over 1000 epochs. The fit generated by these ResNets is acceptable, with a mean absolute error of $1.4161$ and a mean squared error of $4.1011$, based on 1500 points generated within the range $[-3,3]\times[-3,3]\times[-3,3]$. The Figure 7.8 illustrates the sample path of our method, showing early convergence towards to one of the global minimum.
7.3.1 Differences with the Classical Simulated Annealing Algorithm
In this exampleinvolving a function with multiple global minima, we compare the performance of our method with the classical Simulated Annealing (SA) algorithm. The Figure 7.8 illustrates the sample path of our method, showing early convergence towards to one of the global minimum and repetitive behavior in various iterations. In contrast, Figure 7.9 depicts the sample path of the classical SA algorithm. While both methods exhibit comparable behavior in the early iterations, the classical SA algorithm demonstrates significant chaotic exploration and is slower in finding a global minimum. This behavior is evident after iteration $200$, where the function values deviate from convergence.
Figure 7.9: Red line: Sample path of the classical Simulated Annealing algorithm for Multiple Global Minima function. Blue line: fitted trend line.
The Python codes of all the experimental evaluations are available in https://github.com/Nicerova7/output-range-analysis-for-deep-neural-networks-with-simulated-annealing.
@@ 8 Conclusions
This work presents an algorithm based on Simulated Annealing (SA) for output range estimation in Deep Neural Networks (DNNs). This approach effectively adapts to restricted domains, overcoming the limitations posed by high non-linearity and the lack of local geometric information in DNNs. Through rigorous theoretical analysis, we demonstrate the algorithm’s convergence to optimal points, enabling accurate estimation of global minima and maxima.
The experimental evaluation, using complex functions such as Ackley and Drop-Wave, validates the algorithm’s effectiveness, showcasing its ability to handle multiple local minima and multidimensional domains. Compared to classical Simulated Annealing, our method exhibits greater stability and faster convergence. With its simple implementation and versatility, this methodology is applicable to various DNN architectures, providing a practical and reliable tool for neural network verification and validation.